OpenStudy (anonymous):

What value completes the square for the expression? x^2 + 14x

OpenStudy (anonymous):

@surdawi maybe you can help?

OpenStudy (anonymous):

x^2 + 14x = 0 14/2 = 7 ....7^2=49 x^2+ 14x + 49 = 49 (x+7)^2=49 is that what you want

OpenStudy (anonymous):

Yes thanks!